Suppose that the car is stopped from 27m/s within right after the collision and the driver 75kg didn’t push the brake.Seatbelt is non-stretchable and produces a constant force to stop the driver in 10 cm . Calculate the acceleration of the driver and the force that the seatbelt acts on the driver for above two conditions.
1
Expert's answer
2021-09-28T16:54:23-0400

"E=\frac{mv^2}2,"

"A=Fs,"

"E=A,"

"F=\frac{mv^2}{2s}=273~kN,"

"a=\frac Fm=3645~\frac{m}{s^2}."